OpenStudy (anonymous):

Find the slope-intercept form of the line that passes through given the point and has the given slope. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution. (5, 5) m = -3

OpenStudy (anonymous):

times 5x3 see what you get

OpenStudy (mathstudent55):

The slope-intercept form of a line is \(y = mx + b\) where m is the slope, and b is the y-intercept. The y-intercept is the y-coordinate of the point where the line intersects the y-axis. Since you are given the slope, \(m = -3\), you can just replace m with -3 in the equation above: \(y = -3x + b\) Now we need to find b, the y-intercept. We use the equation with the slope, and we use the given point. The given point is (5, 5), so we replace x with 5 and y with 5 and we solve for b. \(5 = -3(5) + b\) \(5 = -15 + b\) Add 15 to both sides: \(10 = b\) \(b = 10\) Now that we know what b is, we replace b with 10. \(y = -3x + b\) \(y = -3x + 10\)